OpenStudy (anonymous):

Solve the proportion 12 over 30 equals x over 5.. What is the value of x? 2 3 4 12

OpenStudy (skullpatrol):

$$\frac{12}{30}=\frac{?}{5}$$

OpenStudy (anonymous):

thats "x"

OpenStudy (anonymous):

Help me solve it please

OpenStudy (skullpatrol):

What do you multiply 30 by to get 5?

OpenStudy (anonymous):

6

OpenStudy (skullpatrol):

$$\frac{1}{6}$$

OpenStudy (skullpatrol):

$$30*\frac{1}{6}=5$$

OpenStudy (skullpatrol):

12 multiplied by $$\frac{1}{6}$$will give you ? or x.

OpenStudy (anonymous):

oh yeah i kinda get it now

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):